40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 168     Entry time: Wed Dec 5 18:08:36 2007
Author: Andrey 
Type: Update 
Category: ASC 
Subject: Optical Lever laser for ETMX is installed 

A new laser with \lambda=633nm has been intalled and the mirror adjusted so that the signal hits the center of the photodetector.

Output power level of that laser is 3.45 +- 0.05 mW.

Only about 0.29mW hits the photodetector.

Cable clips have been used to firmly fix the power supply cable for the laser.

See attached photopicture of the ETMX - "oplev" - optical - table.
Attachment 1: DSC_0199.JPG  367 kB  | Hide | Hide all
DSC_0199.JPG
ELOG V3.1.3-